# Towards Unified Music Emotion Recognition across Dimensional and Categorical Models

### v1

@@ Abstract
One of the most significant challenges in Music Emotion Recognition (MER) comes from the fact that emotion labels can be heterogeneous across datasets with regard to the emotion representation, including categorical (e.g., happy, sad) versus dimensional labels (e.g., valence-arousal). In this paper, we present a unified multitask learning framework that combines these two types of labels and is thus able to be trained on multiple datasets. This framework uses an effective input representation that combines musical features (i.e., key and chords) and MERT embeddings. Moreover, knowledge distillation is employed to transfer the knowledge of teacher models trained on individual datasets to a student model, enhancing its ability to generalize across multiple tasks. To validate our proposed framework, we conducted extensive experiments on a variety of datasets, including MTG-Jamendo, DEAM, PMEmo, and EmoMusic. According to our experimental results, the inclusion of musical features, multitask learning, and knowledge distillation significantly enhances performance. In particular, our model outperforms the state-of-the-art models, including the best-performing model from the MediaEval 2021 competition on the MTG-Jamendo dataset. Our work makes a significant contribution to MER by allowing the combination of categorical and dimensional emotion labels in one unified framework, thus enabling training across datasets.
@@ II-A Single-Task MER Models
The absence of official train/test splits in many datasets, such as the PMEmo dataset [52], makes it difficult to compare models directly, except for the MTG-Jamendo dataset [4], which provides official splits. This enables us to compare our model with those from MediaEval 2021 in our experiments. Despite these achievements, single-task models frequently face challenges in generalizing across various datasets, underscoring the need for research that integrates multiple datasets.
@@ II-C Multitask Learning in MER
Multitask learning (MTL) [8] aims to enhance the performance of multiple related tasks by leveraging shared knowledge among them. Qiu et al. [39] introduced an MTL framework for Symbolic Music Emotion Recognition (SMER). They combines emotion recognition task with auxiliary tasks such as key and velocity classification. The idea is that by forcing the model to learn key and velocity information, it will also better understand the resulting emotion. Evaluated on the EMOPIA [23] and VGMIDI [16] datasets, their approach achieved accuracy improvements of 4.17% and 1.97%, respectively, compared to single-task models. Ji et al. [22] introduced an attention-based deep feature fusion (ADFF) method for valence-arousal prediction within a multitask learning framework. They integrates spatial feature extraction using an adapted VGGNet with a squeeze-and-excitation (SE) attention-based temporal module to enhance affect-salient features. Their model achieved $R^{2}$ scores of 0.4575 for valence and 0.6394 for arousal on the PMEmo dataset [52].
@@ II-E Feature Engineering for MER
Feature engineering is an important facet of Music Emotion Recognition (MER). The earlier works relied on hand-crafted features like MFCCs, chroma features, and rhythmic descriptors [21] which could capture low-level details of the audio, but were not good at depicting higher-level musical semantics [37]. Lately, models like VGGish [19], CLAP [15], and MERT [27], which are trained with self-supervised learning on large-scale datasets, are used to create more elaborate musical embeddings. Among these, MERT has shown remarkable performance in MER. For instance, MERT-95M achieves a PR-AUC of 0.134 and an ROC-AUC of 0.764 on MTG-Jamendo, while the larger MERT-330M model improves these scores to a PR-AUC of 0.14 and an ROC-AUC of 0.765. In addition to embeddings, the high-level musical elements such as key signatures and chord progressions are highly significant in emotion prediction [11]. In this study, the goal is to enhance the MER techniques by combining these large scale embeddings with the high-level musical components.
@@ III-A1 Pre-trained embeddings (MERT)
This augmentation step enhances the robustness of our proposed framework, enabling it to handle a wider range of complex audio data.
@@ III-A2 Musical Features (Chord and Key)
Unlike the segmented and augmented MERT features, musical features (such as chord progressions and key signatures) are calculated over the whole song, representing global harmonic and tonal information. The framework integrates harmonic and tonal information by extracting musical features such as chord progressions and key signatures. A Transformer-based chord recognition model is used for extracting the chord progressions of the song. The chord recognition model achieves Weighted Chord Symbol Recall (WCSR) scores of 83.5% for Root, 80.8% for Thirds, 75.9% for Triads, 71.8% for Sevenths, 65.5% for Tetrads, 82.3% for Maj-min, and 80.8% for MIREX categories [40]. These scores are considered acceptable for this study, as most errors are minor, such as misclassifying A minor as C major.
In our model, each chord is encoded by its root (e.g., D) and quality (e.g., sus4), representing the chord type. After that, these chord progressions are converted into MIDI representations based on music theory. For instance, a C major chord comprises the notes C, E, G, while a C minor 7th chord comprises the notes C, E, G, B. The start and end times of each chord are mapped to define its duration in the MIDI file. This MIDI representation serves as input to a key detection model.
Key detection is performed using the music21 library [12]. To ensure consistency in the representation of chords, we ‘normalize’ them based on the key. Extracted chords from songs in major keys are transposed to C major, while those from minor keys are transposed to A minor. The extracted chord sequences have 13 different chord types such as ‘major’, ‘minor’, ‘diminished’, ‘augmented’, ‘suspended’, and ‘seventh’ chords. The key is encoded based on its mode. For instance, ‘major’ corresponds to C major and ‘minor’ corresponds to A minor.
@@ III-C Knowledge Distillation
where $s$ and $t$ represent the predicted and teacher-generated logits, respectively.
@@ III-D Total Loss Function
The total loss function combines the task-specific loss with the KL divergence loss. For categorical and dimensional datasets, the total loss is:
$L_{\text{Total,Cat}}=\alpha\cdot L_{\text{Cat}}+(1-\alpha)\cdot L_{\text{KD}},$ (10)
$L_{\text{Total,VA}}=\beta\cdot L_{\text{VA}}+(1-\beta)\cdot L_{\text{KD}},$ (11)
respectively, where $\alpha$ and $\beta$ controls the weighting between the task-specific loss and the KL divergence loss.
@@ IV-A Dataset
IV-B Implementation details
To extract MERT features, the audio tracks are segmented into 30-second clips. We leverage the popular MERT-v1-95M model, accessible via Hugging Face [^1] https://huggingface.co/m-a-p/MERT-v1-95M, to obtain the embeddings. Knowledge distillation is used, using teacher models pre-trained on each dataset to guide the student model during multitask learning. For feature extraction, we concatenate the 5th and 6th layer embeddings of MERT and process chord/key features with the Chord Transformer with positional encoding. The hyperparameters for the total loss function (as defined in Equations 10 and 11) are set to $\alpha=0.2$ and $\beta=0.2$. During training, only the relevant loss components are updated based on the dataset type (categorical or dimensional). The models are trained for 200 epochs with a batch size of 8, learning rate of 0.0001 and Adam optimizer. All models are trained on a cluster of four NVIDIA Tesla V100 DGXS GPUs, each with 32 GB of memory. The code is implemented in PyTorch and is available online [^2] https://github.com/AMAAI-Lab/Music2Emotion.
IV-C Performance Metrics
IV-C1 PR-AUC and ROC-AUC
IV-C2 $R^{2}$ Scores for Valence and Arousal
V Results
TABLE II: Comparison of input features and models on performance metrics across datasets.
@@ V Results
TABLE III: Comparison of performance metrics when training on multiple datasets.
@@
In a first experiment, we removed the high-level musical features (i.e., key and chords) from our proposed framework and trained the network using the MERT features only. The performance of the models with these different input feature configurations is shown in Table II. As you can see from Table II, incorporating both the MERT and high-level musical features significantly enhanced the performance of our model. The model with high-level musical features includes the Chord Transformer model. This achieves the best results, including a PR-AUC of 0.1521 and an ROC-AUC of 0.7806 on the MTG-Jamendo dataset. Furthermore, the framework achieves higher $R^{2}_{V}$ and $R^{2}_{A}$ scores for valence and arousal on the dimensional datasets (DEAM, EmoMusic, and PMEmo), demonstrating the importance of including the musical features.
@@
Lastly, we evaluated our proposed model against current state-of-the-art models on the MTG-Jamendo dataset. Table IV shows that our proposed model outperforms all existing models, including the best-performing model, lileonardo [5], from the MediaEval 2021 competition [48] and more recent methods [18, 17, 27]. This substantial improvement makes the proposed framework superior to others in categorical emotion recognition. Since MTG-Jamendo has an official train-validation-test split, our results can be compared directly with prior work. Unfortunately, for other datasets, the lack of such official splits complicates direct comparisons since various studies use different train/test configurations.

### v2

@@ Abstract
One of the most significant challenges in Music Emotion Recognition (MER) comes from the fact that emotion labels can be heterogeneous across datasets with regard to the emotion representation, including categorical (e.g., happy, sad) versus dimensional labels (e.g., valence-arousal). In this paper, we present a unified multitask learning framework that combines these two types of labels and is thus able to be trained on multiple datasets. This framework uses an effective input representation that combines musical features (i.e., key and chords) and MERT embeddings. Moreover, knowledge distillation is employed to transfer the knowledge of teacher models trained on individual datasets to a student model, enhancing its ability to generalize across multiple tasks. To validate our proposed framework, we conducted extensive experiments on a variety of datasets, including MTG-Jamendo, DEAM, PMEmo, and EmoMusic. According to our experimental results, the inclusion of musical features, multitask learning, and knowledge distillation significantly enhances performance. In particular, our model outperforms the state-of-the-art models on the MTG-Jamendo dataset. Our work makes a significant contribution to MER by allowing the combination of categorical and dimensional emotion labels in one unified framework, thus enabling training across datasets.
@@ II-A Single-Task MER Models
The absence of official train/test splits in many datasets, such as the PMEmo dataset [55], makes it difficult to compare models directly, except for the MTG-Jamendo dataset [4], which provides official splits. This enables us to compare our model with those from MediaEval 2021 [5, 40, 48, 34] as well as more recent methods [19, 18, 29] in our experiments. Despite these achievements, single-task models frequently face challenges in generalizing across various datasets, underscoring the need for research that integrates multiple datasets.
@@ II-C Multitask Learning in MER
Multitask learning (MTL) [8] aims to enhance the performance of multiple related tasks by leveraging shared knowledge among them. Qiu et al. [41] introduced an MTL framework for Symbolic Music Emotion Recognition (SMER). They combines emotion recognition task with auxiliary tasks such as key and velocity classification. The idea is that by forcing the model to learn key and velocity information, it will also better understand the resulting emotion. Evaluated on the EMOPIA [24] and VGMIDI [17] datasets, their approach achieved accuracy improvements of 4.17% and 1.97%, respectively, compared to single-task models. Ji et al. [23] introduced an attention-based deep feature fusion (ADFF) method for valence-arousal prediction within a multitask learning framework. They integrate spatial feature extraction using an adapted VGGNet with a squeeze-and-excitation (SE) attention-based temporal module to enhance affect-salient features. Their model achieved $R^{2}$ scores of 0.4575 for valence and 0.6394 for arousal on the PMEmo dataset [55].
@@ II-E Feature Engineering for MER
Feature engineering is a key component in Music Emotion Recognition (MER). Earlier works relied heavily on hand-crafted features such as MFCCs, chroma features, and rhythmic descriptors [22], which capture low-level signal properties but fall short in representing higher-level musical semantics [39]. Other approaches leverage embeddings from audio models trained on large-scale datasets. For example, VGGish [20], a convolutional model trained in a supervised fashion for audio classification, and CLAP [16], which is trained with a contrastive audio-text alignment objective, have been used to extract richer representations of audio content.
Pre-trained encoders have also gained traction in MER due to their ability to learn from unlabeled audio data. Among these, MERT [54] has demonstrated strong performance relative to other pre-trained encoders. For instance, MERT-95M achieves a PR-AUC of 0.134 and an ROC-AUC of 0.764 on the MTG-Jamendo dataset, while the larger MERT-330M model improves these scores slightly to a PR-AUC of 0.14 and an ROC-AUC of 0.765. These results highlight the potential of pre-trained encoders for emotion prediction tasks.
Beyond embeddings, high-level musical features such as key signatures and chord progressions have also been shown to play a significant role in emotion prediction [11]. In this study, we aim to enhance MER by integrating large-scale audio embeddings with these symbolic, high-level musical descriptors.
@@ III-A1 Pre-trained embeddings (MERT)
The intuition behind this augmentation strategy is to expose the model to varying temporal contexts of the same track across training epochs. Emotion in music can vary subtly throughout a piece—different sections might convey slightly different moods due to changes in melody, harmony, rhythm, or instrumentation. By randomly sampling and averaging different combinations of segments, the model learns to generalize across these intra-track variations, rather than overfitting to a single static representation. This not only improves robustness to local variations in musical content but also better reflects the dynamic and evolving nature of emotional expression in music.
@@ III-A2 Musical Features (Chord and Key)
Unlike the segmented and augmented MERT features, musical features (such as chord progressions and key signatures) are calculated over the whole song, representing global harmonic and tonal information. The framework integrates harmonic and tonal information by extracting musical features such as chord progressions and key signatures. A Transformer-based chord recognition model [27] is used for extracting the chord progressions of the song. The chord recognition model was evaluated on the following datasets: a subset of 221 songs from Isophonics [^1] http://isophonics.net/datasets(171 songs by the Beatles, 12 songs by Carole King, 20 songs by Queen and 18 songs by Zweieck), Robbie Williams [14] (65 songs by Robbie Williams), and a subset of 185 songs from UsPop2002222https://github.com/tmc323/Chord-Annotations. The chord recognition model is evaluated using mir_eval metrics [42], achieving Weighted Chord Symbol Recall (WCSR) scores of 83.5% for Root, 80.8% for Thirds, 75.9% for Triads, 71.8% for Sevenths, 65.5% for Tetrads, 82.3% for Maj-min, and 80.8% for the MIREX categories. These scores are considered acceptable for this study, as most errors are minor, such as misclassifying A minor as C major.
In our model, each chord is encoded by its root (e.g., D) and quality (e.g., sus4), representing the chord type, and mapped into an embedding space to capture harmonic patterns. These chord progressions are then converted into MIDI representations based on music theory. For instance, a C major chord comprises the notes C, E, G, while a C minor 7th chord comprises the notes C, E, G, B. The start and end times of each chord are mapped to define its duration in the MIDI file. This MIDI representation serves as input to a key detection model.
Key detection is performed using the music21 library [12]. To ensure consistency in the representation of chords, we ‘normalize’ them based on the key. Extracted chords from songs in major keys are transposed to C major, while those from minor keys are transposed to A minor. The extracted chord sequences have 13 different chord types such as ‘major’, ‘minor’, ‘diminished’, ‘augmented’, ‘suspended’, and ‘seventh’ chords. The key is encoded based on its mode. For instance, ‘major’ corresponds to C major and ‘minor’ corresponds to A minor. These harmonic features—key (as a scalar) and chords (as sequences of embedded tokens)—are integrated with MERT features and used as input to the model.
@@ III-C Knowledge Distillation
where $s$ and $t$ denote the predicted and teacher-generated soft labels, respectively, $c$ is the number of output dimensions (i.e., emotion categories for classification, or 2 for valence and arousal in regression), and $s_{i}$, $t_{i}$ represent the $i$-th elements of the student and teacher distributions, respectively.
@@ III-D Total Loss Function
The total loss function combines the task-specific loss with the KL divergence loss derived from the teacher models. Specifically, for categorical and dimensional datasets, the total loss functions are defined as:
$L_{\text{Total,Cat}}=\alpha\cdot L_{\text{Cat}}+(1-\alpha)\cdot L_{\text{KD,Cat}},$ (10)
$L_{\text{Total,VA}}=\beta\cdot L_{\text{VA}}+(1-\beta)\cdot L_{\text{KD,VA}},$ (11)
where $\alpha$ and $\beta$ are hyperparameters that control the balance between the task-specific loss and the KL divergence loss for classification and regression, respectively. $L_{\text{KD,Cat}}$ and $L_{\text{KD,VA}}$ refer to the KL divergence losses computed using the soft targets generated by the classification and regression teacher models, respectively. $L_{\text{Total,Cat}}$ and $L_{\text{Total,VA}}$ represent the final loss values used to update the classification and regression branches of the multitask model, respectively.
@@ IV-A Dataset
TABLE II: Ablation study on influence of input features on performance metrics across datasets. Here, FNN stands for Feedforward Neural Network.
@@
TABLE III: Comparison of performance metrics for our multitask learning model across multiple (test) datasets, when trained on different training sets. Here, J, D, E, and P stand for the MTG-Jamendo, DEAM, EmoMusic, and PMEmo datasets, respectively.
@@
IV-B Implementation details
To extract MERT features, the audio tracks are segmented into 30-second clips. We leverage the popular MERT-v1-95M model, accessible via Hugging Face [^3] https://huggingface.co/m-a-p/MERT-v1-95M, to obtain the embeddings. Knowledge distillation (KD) is employed, where teacher models pre-trained on each dataset guide the student model during multitask learning. The teacher models are trained separately on each dataset using the same architecture as the student. For feature extraction, we concatenate the 5th and 6th layer embeddings of MERT and process chord/key features with the Chord Transformer with positional encoding. The hyperparameters for the total loss function (as defined in Equations 10 and 11) are set to $\alpha=0.2$ and $\beta=0.2$. During training, only the relevant loss components are updated based on the dataset type (categorical or dimensional). The student and teacher models are trained for 200 epochs with a batch size of 8, a learning rate of 0.0001, and the Adam optimizer. All models, including both the student and teacher models, are trained on a cluster of four NVIDIA Tesla V100 DGXS GPUs, each with 32 GB of memory. The code is implemented in PyTorch and is available online [^4] https://github.com/AMAAI-Lab/Music2Emotion.
IV-C Performance Metrics
IV-C1 PR-AUC and ROC-AUC
IV-C2 $R^{2}$ Scores for Valence and Arousal
IV-D Baseline Models on MTG-Jamendo
To evaluate the effectiveness of our approach, we compare it against several representative models developed for the MTG-Jamendo dataset, as shown in Table IV: 1) lileonardo [5], a deep CNN that applies frequency-dependent convolutions to mel spectrograms, achieving the top score in the MediaEval 2021 competition; 2) SELAB-HCMUS [40], which uses a co-teaching training strategy and short audio segments to enhance performance and reduce training time; 3) Mirable [48], which explores semi-supervised learning through noisy student training and harmonic pitch class profiles (HPCP); 4) UIBK-DBIS [34], an ensemble-based model that clusters emotion labels to train specialized classifiers; 5) Hasumi et al. [19], which proposes classifier group chains to capture tag dependencies for improved music tagging; 6) Greer et al. [18], who introduce M3BERT, a multi-task, self-supervised transformer trained via masked reconstruction and fine-tuned for emotion recognition; and 7) MERT [29], a large-scale acoustic music understanding model (with 95M and 330M parameters) pre-trained using pseudo labels from acoustic and musical teachers, demonstrating strong performance across several music understanding tasks.
V Results
In a first experiment, we removed the high-level musical features (i.e., key and chords) from our proposed framework and trained the network using the MERT features only. The performance of the models with these different input feature configurations is shown in Table II. As can be seen from the table, incorporating both the MERT and high-level musical features significantly enhanced the performance of our model. The model with high-level musical features includes the Chord Transformer model. This achieves the best results, including a PR-AUC of 0.1521 and an ROC-AUC of 0.7806 on the MTG-Jamendo dataset. Furthermore, the framework achieves higher $R^{2}_{V}$ and $R^{2}_{A}$ scores for valence and arousal on the dimensional datasets (DEAM, EmoMusic, and PMEmo), demonstrating the importance of including the musical features.
@@ V Results
Lastly, we evaluated our proposed model on the MTG-Jamendo dataset using its official train-validation-test split, which enables direct comparison with prior work. As shown in Table IV, our model achieves strong performance compared to both earlier baselines such as lileonardo [5] from the MediaEval 2021 competition [50], as well as more recent representative methods [19, 18, 54], which reflect current directions in MER modeling. The substantial improvement over these models suggests that our proposed framework is a promising approach for categorical emotion recognition in music. In contrast, other MER datasets do not offer official splits, which makes direct performance comparison less consistent across studies.